OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (anonymous):

Hey @jim_thompson5910, you busy?

OpenStudy (anonymous):

@sourwing ?

OpenStudy (anonymous):

what do parallel lines have in common?

OpenStudy (anonymous):

the same slope.

OpenStudy (anonymous):

ok, so what is the slope of the given line?

OpenStudy (anonymous):

It says 4..

OpenStudy (anonymous):

no, slope is 4. Solve for y and what do you get?

OpenStudy (anonymous):

\[y = \frac{ 4 }{ 3 }x\]

OpenStudy (anonymous):

it's y = (4/3)x + 4

OpenStudy (anonymous):

I didn't know if you wanted the +4 or not.

OpenStudy (anonymous):

though the 4 is irrelevant, you solved for y incorrectly. anyway, 4/3 is the slope of the new line. Now we just need to find the y-intercept

OpenStudy (anonymous):

y = mx + b, and a point is given

OpenStudy (anonymous):

\[B = ^{-}1\]?

OpenStudy (anonymous):

Reply please?